### v1

@@ Abstract
We propose LEAPS, an algorithm to sample from discrete distributions known up to normalization by learning a rate matrix of a continuous-time Markov chain (CTMC). LEAPS can be seen as a continuous-time formulation of annealed importance sampling and sequential Monte Carlo methods, extended so that the variance of the importance weights is offset by the inclusion of the CTMC. To derive these importance weights, we introduce a set of Radon-Nikodym derivatives of CTMCs over their path measures. Because the computation of these weights is intractable with standard neural network parameterizations of rate matrices, we devise a new compact representation for rate matrices via what we call locally equivariant functions. To parameterize them, we introduce a family of locally equivariant multilayer perceptrons, attention layers, and convolutional networks, and provide an approach to make deep networks that preserve the local equivariance. This property allows us to propose a scalable training algorithm for the rate matrix such that the variance of the importance weights associated to the CTMC are minimal. We demonstrate the efficacy of LEAPS on problems in statistical physics.
@@ 1 Introduction
The most salient approach to such sampling problems is Markov chain Monte Carlo (MCMC) Metropolis et al. (1953); Robert et al. (1999), in which a randomized process is simulated whose equilibrium is the distribution of interest. While powerful and widely applied, MCMC methods can be inefficient as they suffer from slow convergence times into equilibrium, especially for distributions exhibiting multi-modality. Therefore, MCMC is often combined with other techniques that rely on non-equilibrium dynamics, e.g. via annealing from a simpler distribution with annealed importance sampling (AIS) Kahn & Harris (1951); Neal (2001) or sequential Monte Carlo methods (SMC) Doucet et al. (2001). Even then, the variance of these importance weights may be untenably large, and making sampling algorithms more efficient remains an active area of research. Inspired by the rapid progress in deep learning, there has been extensive interest in augmenting contemporary sampling algorithms with learning Noé et al. (2019); Albergo et al. (2019); Gabrié et al. (2022); Nicoli et al. (2020); Matthews et al. (2022), while still maintaining their statistical guarantees.
Recently, there has been rapid progress in development of generative models using techniques from dynamical measure transport, i.e. where data from a base distribution is transformed into samples from the target distribution via flow or diffusion processes Ho et al. (2020); Song et al. (2020); Albergo & Vanden-Eijnden (2022); Albergo et al. (2023); Lipman et al. (2023); Liu et al. (2022). More recently, these models could also be developed for discrete state spaces (Campbell et al., 2022; Gat et al., 2025; Shaul et al., 2024; Campbell et al., 2024) and general state spaces and Markov processes (Holderrieth et al., 2024).
While there have been various developments on adapting these non-equilibrium dynamics for sampling in continuous state spaces Zhang & Chen (2022); Vargas et al. (2023); Máté & Fleuret (2023); Tian et al. (2024); Albergo & Vanden-Eijnden (2024); Richter & Berner (2024); Akhound-Sadegh et al. (2024); Sun et al. (2024), there is a lack of existing literature on such sampling approaches for discrete distributions. However, discrete data are prevalent in various applications, such as in the study of spin models in statistical physics, protein and genomic data, and language. To this end, we provide a new solution to the discrete sampling problem via CTMCs. Our method is similar in spirit to the results in Albergo & Vanden-Eijnden (2024); Vargas et al. (2024) but takes the necessary theoretical and computational leaps to make these approaches possible for discrete distributions. Our main contributions are:
@@ 3 Sampling with CTMCs
where ~$\tilde{\mathbb{P}}[\cdot|X_{t}]$ describes a valid probability distribution for small $h>0$ by the conditions we imposed on $Q_{t}$ (see Equation 4).
@@ 4 Proactive Importance Sampling
Using that $\partial_{t}U_{t}(x)=-\partial_{t}\log\rho_{t}(x)=-\partial_{t}\rho_{t}(x)/\rho_{t}(x)$, we obtain equivalently:
@@
where we defined a new operator $\mathcal{K}_{t}\rho_{t}$. Intuitively, the operator $\mathcal{K}_{t}$ measures the violation from the KFE in log-space and it is intuitive to define $A_{t}$ as the accumulated error of that violation, i.e. as the integral
@@ 5 Proactivate IS via Radon-Nikodym Derivatives
i.e. all trajectories that are continuous from the right with left limits. Such trajectories are commonly called càdlàg trajectories. In other words, jumps (switches between states) happen if and only if $X_{t^{-}}\neq X_{t}$. We consider path distributions (or path measures), i.e. probability distributions over trajectories. For a CTMC $\mathbf{X}=(X_{t})_{0\leq t\leq 1}$ with rate matrix $Q_{t}$ and initial distribution $\mu$, we denote the corresponding path distribution as $\overrightarrow{\mathbb{P}}^{\mu,Q}$ where the arrow $\overrightarrow{\mathbb{P}}$ denotes that we go forward in time. Similarly, we denote with $\overleftarrow{\mathbb{P}}^{\nu,Q^{\prime}}$ a CTMC running in reverse time initialized with $\nu$. We present the following proposition whose proof can be found in Appendix A:
@@
$=\mathbb{E}_{\mathbf{Y}\sim\overrightarrow{\mathbb{P}}^{\rho_{0},Q}}\left[h(X_{t})\frac{\mathrm{d}\overleftarrow{\mathbb{P}}^{\rho_{t},Q^{\prime}}}{\mathrm{d}\overrightarrow{\mathbb{P}}^{\rho_{0},Q}}(\mathbf{X})\right]$
@@
However, to sample efficiently, it is crucial that the IS weights have low variance. Therefore, we will now derive the optimal IS scheme of this form. Ideally the weights will have zero variance - in other words the RND $\frac{\mathrm{d}\overleftarrow{\mathbb{P}}^{\rho_{1},Q^{\prime}}}{\mathrm{d}\overrightarrow{\mathbb{P}}^{\rho_{0},Q}}(\mathbf{X})$ will be constant $=1$. This is the case if and only if the path measures are the same, i.e. if the CTMC in reverse time is a time-reversal of the CTMC running in forward time. It is well-known that this is equivalent to:
@@
A proof can be bound in Appendix B. Note that Theorem 5.2 is useful because we can, in principle, compute $A_{t}$, i.e. there are no unknown variables, and that this holds for arbitrary $Q_{t}$. This theorem can be seen as a discrete state space equivalent of the generalized version of the Jarzynski equality Jarzynski (1997); Vaikuntanathan & Jarzynski (2008) that has also recently been used for sampling in continuous spaces Vargas et al. (2024); Albergo & Vanden-Eijnden (2024). Finally, it is important to note that the IS scheme will not have zero variance if it does not hold that $X_{t}\sim\rho_{t}$.
@@ 6 PINN objective
A proof can be found in Appendix C. Note that we can easily minimize the PINN objective via stochastic gradient descent (see Algorithm 1). It is ”off-policy” as the reference distribution $\mathbb{Q}$ is arbitrary. This objective can be seen as the CTMC equivalent of that in Máté & Fleuret (2023); Albergo & Vanden-Eijnden (2024); Tian et al. (2024); Sun et al. (2024).
@@ 7 Adding Annealed IS and SMC
Algorithm 1 LEAPS training with optional replay buffer
0:  $B$ batch size, $N$ time steps, model $G^{\theta}_{t}$, free energy net $F_{t}^{\phi}$, learning rate $\eta$, replay buffer $\mathcal{B}$.
1:  while not converged do
2:     if use_buffer then
3:        $(X_{t_{m}}^{m},A_{t_{m}}^{m},t_{m})_{m=1,\dots,B}\leftarrow\text{SampleBatch}(\mathcal{B})$
4:     else
5:        $(X_{t_{m}}^{m},A_{t_{m}}^{m},t_{m})_{m=1,\dots,B}\leftarrow$ Algorithm 2
6:     end if
7:     $\mathcal{L}(\theta,\phi)=\frac{1}{B}\sum_{m}|\mathcal{K}_{t_{m}}^{\theta}\rho_{t_{m}}(X_{t_{m}}^{m})-\partial_{t_{m}}F_{t_{m}}^{\phi}|^{2}$
8:     $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}(\theta,\phi)$
9:     $\phi\leftarrow\phi-\eta\,\nabla_{\phi}\mathcal{L}(\theta,\phi)$
10:  end while
@@
1:  Require: $N$ time steps, $M$ walkers, model $F_{t}^{\theta}$, replay buffer $\mathcal{B}$, MCMC kernel $\mathcal{M}_{t}$, density $\rho_{t}$, coeff. $\epsilon_{t}\geq 0$, resample thres. $0\leq\delta\leq 1$
@@ 8 Efficient IS and training via local equivariance
$Q_{t}^{\theta}(y,x)=\begin{cases}0&\text{if }y\notin N(x)\\ Q_{t}^{\theta}(y^{j},j|x)&\text{else if }y_{j}\neq x_{j}\\ -\sum\limits_{i,\tau}Q_{t}^{\theta}(\tau,i|x)&\text{if }x=y\\ \end{cases}$ (19)
This parameterization is commonly used in the context of discrete markov models (”discrete diffusion models”) (Campbell et al., 2022; 2024). With that, the operator $\mathcal{K}_{t}^{\theta}$ in Equation 10 becomes:
@@
where $[z]_{+}=\max(z,0)$ describes the ReLU operation. This representation is not a restriction (see Appendix D for a proof):
@@
Crucially, this representation allows to efficiently compute $\mathcal{K}_{t}^{\theta}$ in one forward pass of the neural network:
@@
It is important to note that with the above construction, we would need to naively evaluate $\rho_{t}(x)$ as often as $d$ times for a single computation of $\mathcal{K}_{t}^{\theta}$. However, note that the sum goes over all neighbors over $x$. Therefore, this can be a considered as computing a discrete gradient. Such ratios can often be computed efficiently, e.g. for many scientific and physical models it is often only 2×2$\times$2 × the computation compared to a single evaluation of $\rho_{t}(x)$.
@@ 9 Design of locally equivariant networks
It remains to be stated how to construct locally equivariant neural networks - a question we turn to in this section. We will focus on three fundamental designs used throughout deep learning: Multilayer perceptrons (MLPs), attention layers, and convolutional neural networks. Usually, tokens are embedded as token vectors $e_{\tau}\in\mathbb{R}^{c_{\text{in}}}$ where $c_{\text{in}}$ is the embedding dimension. We therefore consider the embedded sequence of vectors: $x=(x_{1},\dots,x_{d})\in(\mathbb{R}^{c_{\text{in}}})^{d}$ as the input to the neural network.
@@
where $\sigma(W^{i}x)_{j}$ denotes the $j$-th element of the vector obtained by multiplying the vector $x$ with the matrix $W^{i}$ and applying the activation function $a$ componentwise. One can easily show that this is a locally equivariant neural network corresponding to a MLP with one hidden layer.
Locally-equivariant attention (LEA) layer. Let us consider a self-attention layer operating on keys $k_{j}=k_{j}(x_{j})$, queries $q_{j}=q_{j}(x_{j})$, and values $v_{j}=v_{j}(x_{j})$ each of which is a function of element $x_{j}$. We define the locally equivariant attention layer then as:
@@
Locally-equivariant convolutional (LEC) network. To construct a locally equivariant convolutional neural network (LEC), we assume that our data lies on a grid. A convolutional layer is characterized by a matrix $W\in\mathbb{R}^{(2k-1)\times(2k-1)}$ and its operation is denoted via $k(W)*x$ where $k$ denotes the convolutional kernel with weights $W$. Here, we set the center of $W$ to zero: $W_{kk}=0$ (i.e. such that corresponding location is effectively ignored). To stack such layers, we can make the output of the previous layer feed into the weights of the next layer:
$h_{0}=$ $(1,\dots,1)^{T}$
@@
where $A_{i}\in\mathbb{R}^{d_{i}\times d_{i}},b_{i}\in\mathbb{R}^{d_{i}},c_{i}\in\mathbb{R}^{d_{i}}$ are learnable tensors which operate on each coordinate independently (i.e. a 1x1 convolution) and $\sigma:\mathbb{R}\to\mathbb{R}$ is an activation function to make it non-linear. We call the resulting network
$H_{t}^{\theta}:(\mathbb{R}^{c_{\text{in}}})^{d}\to$ $(\mathbb{R}^{c_{\text{out}}})^{d}$
$x\mapsto$ $(H_{t}^{\theta}(1|x),\dots,H_{t}^{\theta}(d|x))$
the prediction head. Combined with a small network $P_{t}^{\theta}:\mathcal{T}\to\mathbb{R}^{k}$ that we call token projector, we define the full neural network as
In Appendix E, we verify that $G_{t}^{\theta}$ defined in this way is locally equivariant. With this construction, one can stack deep highly complex convolutional neural networks. Note that this convolutional neural network has two (separate) symmetries: it is geometrically translation equivariant and locally equivariant in the sense defined in this work.
Figure 4: Performance metrics of LEAPS on the $L=15,J=0.4,\beta=0.7$ Ising model with the LEA and LEC architectures, using 100 annealing steps for sample generation. Left: Effective sample size of LEAPS samplers over training. Green area denotes an annealing phase during training where $t$ is increased from 00 to $1$. Increasing the depth of LEC significantly improves performance. Center: Difference in the histograms of the magnetization $M(x)$ of configurations as compared to the ground truth set attained from a Glauber dynamics run of 25,000 steps, labeled as $M^{*}$. We denote by ”no transport” the case of using annealed dynamics with just the marginal preserving MCMC updates to show that the transport from $Q_{t}$ is essential in our construction. Right: Comparison of the 2-point correlation function for the LEA and LEC samplers against the Glauber dynamics ground truth.

### v2

@@ Abstract
We propose LEAPS, an algorithm to sample from discrete distributions known up to normalization by learning a rate matrix of a continuous-time Markov chain (CTMC). LEAPS can be seen as a continuous-time formulation of annealed importance sampling and sequential Monte Carlo methods, extended so that the variance of the importance weights is offset by the inclusion of the CTMC. To derive these importance weights, we introduce a set of Radon-Nikodym derivatives of CTMCs over their path measures. Because the computation of these weights is intractable with standard neural network parameterizations of rate matrices, we devise a new compact representation for rate matrices via what we call locally equivariant functions. To parameterize them, we introduce a family of locally equivariant multilayer perceptrons, attention layers, and convolutional networks, and provide an approach to make deep networks that preserve the local equivariance. This property allows us to propose a scalable training algorithm for the rate matrix such that the variance of the importance weights associated to the CTMC are minimal. We demonstrate the efficacy of LEAPS on problems in statistical physics. We provide code in https://github.com/malbergo/leaps/.
Machine Learning, ICML
@@ 1 Introduction
The most salient approach to such sampling problems is Markov chain Monte Carlo (MCMC) (Metropolis et al., 1953; Robert et al., 1999), in which a randomized process is simulated whose equilibrium is the distribution of interest. While powerful and widely applied, MCMC methods can be inefficient as they suffer from slow convergence times into equilibrium, especially for distributions exhibiting multi-modality. Therefore, MCMC is often combined with other techniques that rely on non-equilibrium dynamics, e.g. via annealing from a simpler distribution with annealed importance sampling (AIS) (Kahn & Harris, 1951; Neal, 2001) or sequential Monte Carlo methods (SMC) (Doucet et al., 2001). Even then, the variance of these importance weights may be untenably large, and making sampling algorithms more efficient remains an active area of research. Inspired by the rapid progress in deep learning, there has been extensive interest in augmenting contemporary sampling algorithms with learning (Noé et al., 2019; Albergo et al., 2019; Gabrié et al., 2022; Nicoli et al., 2020; Matthews et al., 2022), while still maintaining their statistical guarantees.
Recently, there has been rapid progress in development of generative models using techniques from dynamical measure transport, i.e. where data from a base distribution is transformed into samples from the target distribution via flow or diffusion processes (Ho et al., 2020; Song et al., 2020; Albergo & Vanden-Eijnden, 2022; Albergo et al., 2023; Lipman et al., 2023; Liu et al., 2022). More recently, these models could also be developed for discrete state spaces (Campbell et al., 2022; Gat et al., 2025; Shaul et al., 2024; Campbell et al., 2024) and general state spaces and Markov processes (Holderrieth et al., 2024).
While there have been various developments on adapting these non-equilibrium dynamics for sampling in continuous state spaces (Zhang & Chen, 2022; Vargas et al., 2023; Máté & Fleuret, 2023; Tian et al., 2024; Albergo & Vanden-Eijnden, 2024; Richter & Berner, 2024; Akhound-Sadegh et al., 2024; Sun et al., 2024), there is a lack of existing literature on such sampling approaches for discrete distributions. However, discrete data are prevalent in various applications, such as in the study of spin models in statistical physics, protein and genomic data, and language. To this end, we provide a new solution to the discrete sampling problem via CTMCs. Our method is similar in spirit to the results in (Albergo & Vanden-Eijnden, 2024; Vargas et al., 2024) but takes the necessary theoretical and computational leaps to make these approaches possible for discrete distributions. Our main contributions are:
@@ 3 Sampling with CTMCs
where $\tilde{\mathbb{P}}[\cdot|X_{t}]$ describes a valid probability distribution for small $h>0$ by the conditions we imposed on $Q_{t}$ (see Equation 4).
@@ 4 Proactive Importance Sampling
Using that $\partial_{t}\rho_{t}(x)/\rho_{t}(x)=\partial_{t}\log\rho_{t}(x)=\partial_{t}F_{t}-\partial_{t}U_{t}(x)$, we obtain after dropping the unknown time-dependent constant $\partial_{t}F_{t}$:
@@
where we defined a new operator $\mathcal{K}_{t}\rho_{t}$. Intuitively, the operator $\mathcal{K}_{t}$ measures the violation from the KFE in log-space (up to $F_{t}$) and it is intuitive to define $A_{t}$ as the accumulated error of that violation, i.e. as the integral
@@ 5 Proactivate IS via Radon-Nikodym Derivatives
[…]
@@
$=\mathbb{E}_{\mathbf{X}\sim\overrightarrow{\mathbb{P}}^{\rho_{0},Q}}\left[h(X_{t})\frac{\mathrm{d}\overleftarrow{\mathbb{P}}^{\rho_{t},Q^{\prime}}}{\mathrm{d}\overrightarrow{\mathbb{P}}^{\rho_{0},Q}}(\mathbf{X})\right]$
@@
However, to sample efficiently, it is crucial that the IS weights have low variance. Therefore, we will now derive the optimal IS scheme of this form. Ideally the weights will have zero variance - in other words the RND $\frac{\mathrm{d}\overleftarrow{\mathbb{P}}^{\rho_{1},Q^{\prime}}}{\mathrm{d}\overrightarrow{\mathbb{P}}^{\rho_{0},Q}}(\mathbf{X})$ will be constant =1=1. This is the case if and only if the path measures are the same, i.e. if the CTMC in reverse time is a time-reversal of the CTMC running in forward time. It is well-known that this is equivalent to:
@@
A proof can be bound in Appendix B. Note that Theorem 5.2 is useful because we can, in principle, compute $A_{t}$, i.e. there are no unknown variables, and that this holds for arbitrary $Q_{t}$. This theorem can be seen as a discrete state space equivalent of the generalized version of the Jarzynski equality (Jarzynski, 1997; Vaikuntanathan & Jarzynski, 2008) that has also recently been used for sampling in continuous spaces (Vargas et al., 2024; Albergo & Vanden-Eijnden, 2024). Finally, it is important to note that the IS scheme will not have zero variance if it does not hold that $X_{t}\sim\rho_{t}$.
@@ 6 PINN Objective
A proof can be found in Appendix C. Note that we can easily minimize the PINN objective via stochastic gradient descent (see Algorithm 2). It is ”off-policy” as the reference distribution $\mathbb{Q}$ is arbitrary. This objective can be seen as the CTMC equivalent of that in (Máté & Fleuret, 2023; Albergo & Vanden-Eijnden, 2024; Tian et al., 2024; Sun et al., 2024).
@@ 7 Adding Annealed IS and SMC
1:Require: NN time steps, $M$ walkers, model $F_{t}^{\theta}$, replay buffer $\mathcal{B}$, MCMC kernel $\mathcal{M}_{t}$, density $\rho_{t}$, coeff. $\epsilon_{t}\geq 0$, resample thres. $0\leq\delta\leq 1$
@@ 8 Efficient IS and Training via Local Equivariance
[…]
This parameterization is commonly used in the context of discrete markov models (”discrete diffusion models”) (Campbell et al., 2022, 2024). With that, the operator $\mathcal{K}_{t}^{\theta}$ in Equation 10 becomes:
@@
where $z]_{+}=\max(z,0)$ describes the ReLU operation. This representation is not a restriction (see Appendix D for a proof):
@@
Note that this representation implies that rates only go from $x$ to $y$ or from $y$ to $x$, and in that sense are kinetically favorable (Shaul et al., 2024). Crucially, this representation allows to efficiently compute $\mathcal{K}_{t}^{\theta}$ in one forward pass of the neural network:
@@
It is important to note that with the above construction, we would need to naively evaluate $\rho_{t}(x)$ as often as $d$ times for a single computation of $\mathcal{K}_{t}^{\theta}$. However, note that the sum goes over all neighbors over $x$. Therefore, this can be a considered as computing a discrete gradient. Such ratios can often be computed efficiently, e.g. for many scientific and physical models it is often only $2\times$ the computation compared to a single evaluation of $\rho_{t}(x)$.
@@ 9 Design of Locally Equivariant Networks
It remains to be stated how to construct locally equivariant neural networks - a question we turn to in this section. We will focus on three fundamental designs used throughout deep learning: Multilayer perceptrons (MLPs), attention layers, and convolutional neural networks. Usually, tokens are embedded as token vectors $e_{\tau}\in\mathbb{R}^{c_{\text{in}}}$ where $c_{\text{in}}$ is the embedding dimension. We therefore consider the embedded sequence of vectors: $x=(x_{1},\dots,x_{d})\in(\mathbb{R}^{c_{\text{in}}})^{d}$ as the input to the neural network. The specific designs we discuss here are based on two components: (1) a locally-invariant prediction head given by
$H_{t}^{\theta}:(\mathbb{R}^{c_{\text{in}}})^{d}\to$ $(\mathbb{R}^{c_{\text{out}}})^{d}$
$x\mapsto$ $(H_{t}^{\theta}(1|x),\dots,H_{t}^{\theta}(d|x))$
i.e. $H_{t}^{\theta}$ fulfills that $H_{t}^{\theta}(i|\text{Swap}(x,i,\tau))=H_{t}^{\theta}(i|x)$ for all $x,i,\tau$. (2) This is combined with a small network $P_{t}^{\theta}:\mathcal{T}\to\mathbb{R}^{k}$ that we call token projector. The full network is then given by
In Appendix E, we verify that $G_{t}^{\theta}$ defined in this way is locally equivariant. We discuss specific instances of this design.
@@
where $\sigma(W^{i}x)_{j}$ denotes the $j$-th element of the vector obtained by multiplying the vector $x$ with the matrix $W^{i}$ and applying the activation function $a$ componentwise. This is a locally equivariant MLP with one hidden layer.
Locally-equivariant attention (LEA) layer. Let us consider a self-attention layer operating on keys $k_{j}=k_{j}(x_{j})$, queries $q_{j}=q_{j}(x_{j})$, and values $v_{j}=v_{j}(x_{j})$ - each of which is a function of element $x_{j}$. We define the locally equivariant attention layer then as:
@@
Locally-equivariant convolutional (LEC) network. To construct a locally equivariant convolutional neural network (LEC), we assume that our data lies on a grid. A convolutional layer is characterized by a matrix $W\in\mathbb{R}^{(2k-1)\times(2k-1)}$ and its operation is denoted via k(W)xk(W)*x where $k$ denotes the convolutional kernel with weights $W$. Here, we set the center of $W$ to zero: $W_{kk}=0$ (i.e. such that corresponding location is effectively ignored). To stack such layers, we can make the output of the previous layer feed into the weights of the next layer:
@@
where $A_{i}\in\mathbb{R}^{d_{i}\times d_{i}},b_{i}\in\mathbb{R}^{d_{i}},c_{i}\in\mathbb{R}^{d_{i}}$ are learnable tensors which operate on each coordinate independently (i.e. a 1x1 convolution) and $\sigma:\mathbb{R}\to\mathbb{R}$ is an activation function to make it non-linear. With this construction, one can stack deep highly complex convolutional neural networks. Note that this convolutional neural network has two (separate) symmetries: it is geometrically translation equivariant and locally equivariant in the sense defined in this work.